OpenStudy (anonymous):

Which of the following polynomials does not have real roots? A. -x2 + 4x + 6 B. x2 - 17x - 7 C. x2 + 4x + 8 D. 2x2 - 13

OpenStudy (anonymous):

is it D ?

zepdrix (zepdrix):

I guess for each of these, you should just worry about the `descriminant`. When given: \(\large\rm ax^2+bx+c\) Solutions of x are given by:\[\large\rm x=\frac{-b\pm\sqrt{\color{orangered}{b^2-4ac}}}{2a}\]We really only care about this orange part, when it's negative, we have no real solutions.

zepdrix (zepdrix):

Let's look at D a sec, It has this form:\[\large\rm 2x^2+0x-13\]No x's, so I put a placeholder to show what my b coefficient is.

zepdrix (zepdrix):

\[\large\rm b^2-4ac\quad=0^2-4(2)(-13)\]D turns out to be a `positive` number, ya?

OpenStudy (anonymous):

so it would be B

OpenStudy (anonymous):

or A

OpenStudy (anonymous):

because you said it would have to be positive

zepdrix (zepdrix):

Option A:\[\large\rm -1x^2+4x+6\]gives,\[\large\rm b^2-4ac\quad=(-1)^2-4(4)(6)\] Option B:\[\large\rm 1x^2 - 17x - 7\]gives,\[\large\rm b^2-4ac\quad=(-17)^2-4(1)(-7)\]

zepdrix (zepdrix):

Yes, it has to be negative result to give us the answer we're looking for. Which one is that? :)

zepdrix (zepdrix):

positive discriminant = real roots We're looking for this situation: negative discriminant = no real roots

OpenStudy (anonymous):

so is it A ? because gives me a positive answer

zepdrix (zepdrix):

A gives you a positive? Are you sure?\[\large\rm b^2-4ac\quad=(-1)^2-4(4)(6)\quad= 1-96\quad= -95\]

OpenStudy (anonymous):

i meant to put b for the positive answer

zepdrix (zepdrix):

We were looking for the `negative` though, so yes option A :)

OpenStudy (anonymous):

thank you , (: